Booth encoding for -52 in 8-bit is
a) 0-1011-100
b) 0-1011-101
c) 0-101-1001
d) 0-1010-100
1
Expert's answer
2012-10-02T07:51:29-0400
0-1010-100 = -64+16-4 = -52
answer is d